# MNSC-0001-2024.00

### 15 Mar 2026 v1

@@ 8 Proofs of Theoretical Results in Section 5
by (67), (68) and (72). Therefore, we have ${\bf c}^{0}\in{\cal B}(y,\zeta^{0})$. Consequently, since $\Phi^{\bf A}_{i}(y,{\bf x}_{-i}^{0})=e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c}^{0})$ and $\tilde{g}_{i}(y)\leq e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c}^{0})$, $\tilde{g}_{i}(y)\leq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})$ holds.
@@
Hence, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq e_{i}+c_{i}^{1}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c}^{1})$ for any feasible $\zeta^{1}$ and ${\bf c}^{1}$, which implies that $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq\tilde{g}_{i}(y)$. As a consequence, (71) follows.
@@
$\tilde{\zeta}\in\mathop{\arg\min}_{\zeta\in{\cal Z}\left(\bar{x}^{\bf A}_{i}/2\right)}\left\{\min_{{\bf c}\in{\cal B}\left(\bar{x}^{\bf A}_{i}/2,\zeta\right)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c})\right\}\right\},$ (73)
@@
since $\zeta^{*}\leq\tilde{\zeta}$ and $Q(x)$ is decreasing in $x$ by Assumption 4.1(ii). Further, we claim that
@@
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\bm{\pi}(\tilde{{\bf c}})>d_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf d}).$ (93)
To that end, we use the following lemma whose proof is deferred to Appendix 9.
Lemma 8.5
@@
If $\bm{\pi}(\tilde{{\bf c}})=\bm{\pi}({\bf d})$, (93) holds by Lemma 8.5. Otherwise, let $\hat{{\bf c}}\coloneqq{\bf d}\wedge\tilde{{\bf c}}$. Then, $\hat{c}_{l}=\tilde{c}_{l}$ by (84), $\hat{c}_{i}=d_{i}$ by Lemma 8.5, and for all $j\neq i,l$,
@@
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\bm{\pi}(\tilde{{\bf c}})-\left(d_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf d})\right)$
$=\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\bm{\pi}(\tilde{{\bf c}})-\bm{\pi}({\bf d})\right)$
$\geq\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\bm{\pi}(\hat{{\bf c}})-\bm{\pi}({\bf d})\right)$ (95)
$>\tilde{c}_{i}-d_{i}-\left\|\bm{\pi}(\hat{{\bf c}})-\bm{\pi}({\bf d})\right\|_{1}$ (96)
@@
where (95) holds since ${\bf A}_{i}$ is nonnegative and $\bm{\pi}(\cdot)$ is increasing (see EisenbergNoe2001, Lemma 5), (96) is satisfied since $A_{ki}\leq\sum_{j=1}^{m}A_{kj}<1$ for all $k\in[m]$, (97) stems from the fact that $\bm{\pi}(\cdot)$ is $\ell_{1}$-nonexpansive (see EisenbergNoe2001, Lemma 5), and (98) follows since we have ${\bf 1}^{\top}\tilde{{\bf c}}=\tilde{\eta}$ and ${\bf 1}^{\top}{\bf d}=\eta^{*}$. Consequently, by (80) and Lemma 8.5, (93) follows.
@@
$>e_{i}+d_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf d})$
$\geq\min_{\zeta\in{\cal Z}(x)}\min_{{\bf c}\in{\cal B}(x,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\bm{\pi}({\bf c})\right\}$ (100)
@@
[…]
9 Proofs for Auxiliary Results
Proof 9.1
Proof of Lemma 8.2. Assume by contradiction that $\bar{p}^{\bf A}>p_{2}^{\bf A}$. For $k=0,1,2,\ldots,$ define
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\bm{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)~~\text{and}~~\bm{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\bm{\lambda}_{k}^{\dagger}\right),$ (102)
@@ 9 Proofs for Auxiliary Results
$\rho_{0}^{\dagger}\coloneqq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)~~\text{and}~~\bm{\lambda}_{0}^{\dagger}\coloneqq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\bm{\pi}^{\bf A}\right).$ (103)
Fix ${\bf x}\leq\bar{{\bf x}}^{\bf A}$. Then, $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}>p_{2}^{\bf A}$. Moreover, by (15) and Assumption 4.1(ii), we have
$p_{2}^{\bf A}$ $=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)$
$\geq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)$
@@
Consequently, we observe that $\bm{\pi}^{\bf A}={\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\bm{\pi}^{\bf A}\right)\geq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\bm{\pi}^{\bf A}\right)=\bm{\lambda}_{0}^{\dagger}$,
If $\rho_{k}^{\dagger}\leq p_{2}^{\bf A}$ and $\bm{\lambda}_{k}^{\dagger}\leq\bm{\pi}^{\bf A}$ for some $k\geq 0$, then
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\bm{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)\leq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q({\bf 1}^{\top}{\bf x})-{\bf A}^{\top}\bm{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)=p_{2}^{\bf A}$ (104)
@@
$\bm{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\bm{\lambda}_{k}^{\dagger}\right)\leq{\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\bm{\pi}^{\bf A}\right)=\bm{\pi}^{\bf A}.$ (105)
[…]
Proof 9.2
Proof of Lemma 8.5. We first let $\delta=\tilde{\zeta}-\zeta^{*}$. Then, by (76),
@@
[…]

### 17 Mar 2026 v2

@@ Appendix B Proofs of Theoretical Results in Section 5
by (67), (68) and (72). Therefore, we have ${\bf c}^{0}\in{\cal B}(y,\zeta^{0})$. Consequently, since $\Phi^{\bf A}_{i}(y,{\bf x}_{-i}^{0})=e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{0})$ and $\tilde{g}_{i}(y)\leq e_{i}+c_{i}^{0}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{0})$, $\tilde{g}_{i}(y)\leq\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})$ holds.
@@
Hence, $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq e_{i}+c_{i}^{1}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c}^{1})$ for any feasible $\zeta^{1}$ and ${\bf c}^{1}$, which implies that $\min_{{\bf x}_{-i}\in[{\bf 0},\bar{{\bf x}}^{\bf A}_{-i}]}\Phi^{\bf A}_{i}(y,{\bf x}_{-i})\leq\tilde{g}_{i}(y)$. As a consequence, (71) follows.
@@
$\tilde{\zeta}\in\mathop{\arg\min}_{\zeta\in{\cal Z}\left(\bar{x}^{\bf A}_{i}/2\right)}\left\{\min_{{\bf c}\in{\cal B}\left(\bar{x}^{\bf A}_{i}/2,\zeta\right)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}\right\},$ (73)
@@
since $\zeta^{*}\leq\tilde{\zeta}$ and $Q(x)$ is decreasing in $x$ by Assumption 1(ii). Further, we claim that
@@
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}(\tilde{{\bf c}})>d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d}).$ (93)
To that end, we use the following lemma whose proof is deferred to Appendix C.
Lemma 4.
@@
If $\boldsymbol{\pi}(\tilde{{\bf c}})=\boldsymbol{\pi}({\bf d})$, (93) holds by Lemma 4. Otherwise, let $\hat{{\bf c}}\coloneqq{\bf d}\wedge\tilde{{\bf c}}$. Then, $\hat{c}_{l}=\tilde{c}_{l}$ by (84), $\hat{c}_{i}=d_{i}$ by Lemma 4, and for all $j\neq i,l$,
@@
$\tilde{c}_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}(\tilde{{\bf c}})-\left(d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d})\right)$
$=\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\boldsymbol{\pi}(\tilde{{\bf c}})-\boldsymbol{\pi}({\bf d})\right)$
$\geq\tilde{c}_{i}-d_{i}+{\bf A}_{i}^{\top}\left(\boldsymbol{\pi}(\hat{{\bf c}})-\boldsymbol{\pi}({\bf d})\right)$ (95)
$>\tilde{c}_{i}-d_{i}-\left\|\boldsymbol{\pi}(\hat{{\bf c}})-\boldsymbol{\pi}({\bf d})\right\|_{1}$ (96)
@@
where (95) holds since ${\bf A}_{i}$ is nonnegative and $\boldsymbol{\pi}(\cdot)$ is increasing (see Eisenberg and Noe, 2001, Lemma 5), (96) is satisfied since $A_{ki}\leq\sum_{j=1}^{m}A_{kj}<1$ for all $k\in[m]$, (97) stems from the fact that $\boldsymbol{\pi}(\cdot)$ is $\ell_{1}$-nonexpansive (see Eisenberg and Noe, 2001, Lemma 5), and (98) follows since we have ${\bf 1}^{\top}\tilde{{\bf c}}=\tilde{\eta}$ and ${\bf 1}^{\top}{\bf d}=\eta^{*}$. Consequently, by (80) and Lemma 4, (93) follows.
@@
$>e_{i}+d_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf d})$
$\geq\min_{\zeta\in{\cal Z}(x)}\min_{{\bf c}\in{\cal B}(x,\zeta)}\left\{e_{i}+c_{i}+{\bf A}_{i}^{\top}\boldsymbol{\pi}({\bf c})\right\}$ (100)
@@
where (99) stems from (71), (73) and (74), and (100) holds since $\zeta^{*}\in{\cal Z}(x)$ and ${\bf d}\in{\cal B}(x,\zeta^{*})$. Then, by (71), we have (66), and the proof is complete. ∎
Appendix C Proofs for Auxiliary Results
Proof.
Proof of Lemma 3. Assume by contradiction that $\bar{p}^{\bf A}>p_{2}^{\bf A}$. For $k=0,1,2,\ldots,$ define
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)~~\text{and}~~\boldsymbol{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right),$ (102)
@@ Appendix C Proofs for Auxiliary Results
$\rho_{0}^{\dagger}\coloneqq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)~~\text{and}~~\boldsymbol{\lambda}_{0}^{\dagger}\coloneqq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right).$ (103)
Fix ${\bf x}\leq\bar{{\bf x}}^{\bf A}$. Then, $p_{1}\coloneqq Q({\bf 1}^{\top}{\bf x})\geq Q({\bf 1}^{\top}\bar{{\bf x}}^{\bf A})=\bar{p}^{\bf A}>p_{2}^{\bf A}$. Moreover, by (15) and Assumption 1(ii), we have
$p_{2}^{\bf A}$ $=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}p_{1}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)$
$\geq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge z_{j}\right)$
@@
Consequently, we observe that $\boldsymbol{\pi}^{\bf A}={\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)\geq{\bf L}\wedge\left({\bf e}+{\bf z}\rho_{0}^{\dagger}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)=\boldsymbol{\lambda}_{0}^{\dagger}$,
If $\rho_{k}^{\dagger}\leq p_{2}^{\bf A}$ and $\boldsymbol{\lambda}_{k}^{\dagger}\leq\boldsymbol{\pi}^{\bf A}$ for some $k\geq 0$, then
$\rho_{k+1}^{\dagger}=Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)^{+}}{\rho_{k}^{\dagger}}\wedge z_{j}\right)\leq Q\left(\sum_{j=1}^{m}\frac{\left(L_{j}-e_{j}-x_{j}Q({\bf 1}^{\top}{\bf x})-{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)^{+}}{p_{2}^{\bf A}}\wedge\left(z_{j}-x_{j}\right)\right)=p_{2}^{\bf A}$ (104)
@@
$\boldsymbol{\lambda}_{k+1}^{\dagger}={\bf L}\wedge\left({\bf e}+{\bf z}\rho_{k}^{\dagger}+{\bf A}^{\top}\boldsymbol{\lambda}_{k}^{\dagger}\right)\leq{\bf L}\wedge\left({\bf e}+{\bf x}p_{1}+({\bf z}-{\bf x})p_{2}^{\bf A}+{\bf A}^{\top}\boldsymbol{\pi}^{\bf A}\right)=\boldsymbol{\pi}^{\bf A}.$ (105)
Hence, by induction, we obtain $\rho_{k}^{\dagger}\leq p_{2}^{\bf A}$ and $\boldsymbol{\lambda}_{k}\leq\boldsymbol{\pi}^{\bf A}$ for all $k$, suggesting that $\bar{p}^{\bf A}\leq p_{2}^{\bf A}$ and $\bar{\boldsymbol{\pi}}^{\bf A}\leq\boldsymbol{\pi}^{\bf A}$ since $\rho_{k}^{\dagger}\rightarrow\bar{p}^{\bf A}$ and $\boldsymbol{\lambda}_{k}^{\dagger}\rightarrow\bar{\boldsymbol{\pi}}^{\bf A}$ as $k$ increases. This contradicts the assumption and indicates that $\bar{p}^{\bf A}\leq p_{2}^{\bf A}$. ∎
Proof.
Proof of Lemma 4. We first let $\delta=\tilde{\zeta}-\zeta^{*}$. Then, by (76),
@@
which implies that the last equality of (108) holds. This completes the proof. ∎